Q1. You purchase a house for $750,000, you are able to make a down payment constituting 1/3 of the cost of the house and take a mortgage to cover the rest. The mortgage you negotiate with the bank is a 30 year, 5% mortgage compounded semi-annually, and you make monthly mortgage payments. a) Under these terms, what is your monthly mortgage payment? b) Assuming that you can only afford monthly payments of $2500. Given the same rate above (5% mortgage with semi-annual compounding), how much will you still owe the bank at the end of the 30 year period? Q2. Your uncle finds out that you are taking this course and asks you to help him plan for retirement. Your uncle turns 55 today and plans to retire in 10 years when he is 65. His current savings are $2 Million dollars and he expects to contribute $100,000 annually to his retirement account for the next 10 years. With the first payment taking place one year from today, and the last payment taking place in 10 years. After retirement, your uncle hopes that his retirement account will last him 20 years; he desires to consume $300,000 starting one year after retirement and this amount will grow by 4% per year as he gets older. Assume that the market interest rate is 6% per year for the entire planning horizon. a) How much will your uncle have in his account when he retires in 10 years? b) Will your uncle have enough money in his retirement account to meet his spending goals after retirement? Please support your answer
Answer #1

Question 1:

(a) Down payment = 1/3*750,000 = 250,000

Loan amount = 750,000-250,000 = 500,000

Effective annual interest (1+0.05/2)^2 -1 = 5.0625% = 0.050625

Monthly payment is calculated as =PMT(rate,nper,pv) in excel where rate = 0.050625/12, nper = 30*12 = 360, pv = 500,000

Monthly payment =PMT(0.050625/12,360,500000) = 2,703.24

Monthly mortgage payment = $2,703.24

(b) Amount you still owe the bank = (2703.24-2500)*360 = $73,166.11

Amount you still owe the bank =$73,166.11

Note: We have answered one full question with all sub-parts. Only one full question will be answered at a time. Kindly post the other question separately for experts to answer
Answer #2

 

1.       Jimmy hopes that the money from his retirement funds plus what he makes on the sale of his current house when he is 65, will be enough to allow him to buy the retirement home. Can this goal be realized?                             

 

 


source: Business
answered by: Sarah Brooks